OpenStudy (anonymous):

Completely factor the polynomial 4x^3+16x^2+4x+16

OpenStudy (cwrw238):

can you see the GCF of the 4 terms?

OpenStudy (anonymous):

4x?

OpenStudy (cwrw238):

no - because of the 16 GCF = 4

OpenStudy (anonymous):

ok so that drops off the x because no x is in that term

OpenStudy (cwrw238):

so we get 4(x^3 + 4x^2 + x + 4)

OpenStudy (cwrw238):

now you need to factor the parentheses as last term is 4 x+ 2 or x+ 4 are possible factors if x+ 2 is a factor then f(-2) = 0

OpenStudy (anonymous):

OK that part confuses me how does it change the last part of it

OpenStudy (cwrw238):

i'm applying the factor theorem - if x-a is a factor of f(x) then f(a) = 0

OpenStudy (cwrw238):

if (x+a) is a factor then f(-a) = 0

OpenStudy (cwrw238):

i 'll try to see if (x+4) is a factor of the polynomial we have : f(-4) = (-4)^3 + 4(-4)^2 - 4 + 4 = -64 + 64 - 4 + 4 = 0

OpenStudy (cwrw238):

that means (x + 4) is one of the factors

OpenStudy (cwrw238):

to find the other factors we need to divide the polynomial by x+4